OpenStudy (anonymous):

The limit of cosx tanx /x as x approaches 0

OpenStudy (anonymous):

does it help to know that \[\cos(x)\tan(x)=\cos(x)\frac{\sin(x)}{\cos(x)}=\sin(x)\]?

OpenStudy (anonymous):

Yeah a bit

OpenStudy (anonymous):

it should because this is really \[\lim_{x\to 0}\frac{\sin(x)}{x}\] in disguise

OpenStudy (anonymous):

sin(x)/cos(x) = sin (x) or sin(x)/cos(x) = tan(x)

OpenStudy (anonymous):

??